Calculating the momentum of a 2.0g feather moving at 3m/s?

Then the kinetic energy of the feather from problem 1 in joules
1
Expert's answer
2020-04-10T08:53:16-0400

The momentum of a feather

"p=mv=(0.002\:\rm kg)(3.0\:\rm m/s)=0.006 \:\rm kg\:m/s."

The kinetic energy of a feather

"K=\frac{mv^2}{2}=\frac{(0.002\:\rm kg)(3.0\:\rm m/s)^2}{2}=0.009 \:\rm J."